OpenStudy (anonymous):

How many different samples of size n can be chosen from a finite population of size a) N=7 b) N=24?

OpenStudy (anonymous):

sorry, size n=2

OpenStudy (anonymous):

\[\dbinom{7}{2}=\frac{7\times 6}{2}=21\]\]

OpenStudy (anonymous):

similarly \[\dbinom{24}{2}=\frac{24\times 23}{2}\]

OpenStudy (anonymous):

awesome, thanks